OpenStudy (anonymous):

what is the solution to 2x^5=-64

OpenStudy (sgadi):

x=-2

OpenStudy (anonymous):

how did you get that

OpenStudy (sgadi):

x^5=-32 by factorizing 32, we get x^5=(-2)^5 so x=-2

OpenStudy (anonymous):

ok